OpenStudy (anonymous):

Sole the equation 9x+5=-4+5x+25

OpenStudy (adunb8):

are you solving for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well lets arrange the equation.. move all the x terms to one side (left for example) and the others to the right

OpenStudy (anonymous):

do you think you can do it ?

OpenStudy (anonymous):

sorry about the spelling but yes

OpenStudy (adunb8):

move all the x to the one side and all the number to the other side then add,subtract and divide and volia!

OpenStudy (anonymous):

no? ive been stuck on equations for awhile

OpenStudy (anonymous):

ok.. if you want to move "5x" to the left side its the same as subtracting 5x from both sides of the equation : 9x + 5 - 5x = -4 + 25

OpenStudy (anonymous):

now .. we still got that +5 on the left side .. so in the same manner we can move it to the right side .. try to move it to the right side by yourself

OpenStudy (anonymous):

9x+5-5x=-4+25-5x Like this right?

OpenStudy (anonymous):

no, look when i subtracted 5x from both sides i left with : 9x + 5 - 5x = -4 + 25 ill show you the process: this is the original : 9x+5=-4+5x+25 now : 9x + 5 -5x = -4 +5x +25 -5x 9x + 5 - 5x = -4 +25

OpenStudy (anonymous):

do you understand ?

OpenStudy (anonymous):

kind of. . can you do the whole thing step by step just so i can see thw whole process please?

OpenStudy (anonymous):

yes 9x+5=-4+5x+25 subtract 5x from both sides 9x + 5 -5x = -4 +5x +25 -5x 9x + 5 - 5x = -4 +25 subtract 5 from both sides: 9x + 5 - 5x -5 = -4 +25 -5 9x- 5x = -4 +25 -5 4x = 16 divide both sides by 4: x = 4

OpenStudy (anonymous):

check yourself : plug x=4 in your original equation to check if it is really the solution : left side : 9*4 + 5 = 41 right side: -4 + 5*4 +25 = 41 41 = 41 its good.

OpenStudy (anonymous):

thanks you thst helped!!

OpenStudy (anonymous):

yw:)